a man wishes to swim across a river 0.5 km wide. If he can swim at the rate 2 km/h in still water and the river flows at the rate of 1km/h. the angle (w.r.t. the flow of the river) along which he should swim so as to reach a point exactly opposite his starting point, should be :a) 60°b) 120°c) 145°d) 90°

Grade:12th pass

1 Answers

Aman gohel
6 years ago
V(m)=2Km/hrV(r)=1Km/hrAs he has to reach just opposite he has to swim leftwards at an angle and river flows rightwards,=>v(m/r)=√(2^2-1^2)=√(4-1)=√3cosX=√3/2=>X=30°(Ans)